OpenStudy (anonymous):

A country's population in 1995 was 228 million. In 2001 it was 230 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (dumbcow):

\[P _{t}=P _{0}e ^{kt}\] Find k \[230 = 228*e ^{6k}\] \[k = \frac{\ln 230/228}{6}\] Substitute this value in to find P15 2010-1995 = 15 \[P _{15}=228*e ^{15*\ln (230/228)/6}\]

OpenStudy (anonymous):

can you solve it for me? i'm not sure how to

OpenStudy (dumbcow):

i did solve it just calculate P15 using a calculator

OpenStudy (anonymous):

how do i do that?

OpenStudy (dumbcow):

ok can you find ln(230/228) ?

OpenStudy (anonymous):

not sure

OpenStudy (dumbcow):

press the ln button and then "(230/228)" did you understand any of the steps i did?

OpenStudy (anonymous):

i got 0.008733 from the In 230/228. then what??????????????????

OpenStudy (anonymous):

They Answer to that is 232. Youre welcome